# Expressivity of Neural Networks with Random Weights and Learned Biases

### v1

@@ 3.3.2 Non-autonomous dynamical systems
We found that both the fully-trained and bias-only networks accurately predicted future points of the system, evidenced by a consistent $R^{2}$ metric of >0.99 ($n$=5) on a window of the generated Lorenz attractor held out from training (Fig. 4A). Furthermore, the models showed remarkable stability when predicting windows that were several times larger than the ones used during training, continuing to reconstruct the system without a notable change in accuracy (Fig. 4B). However, when the networks were fed their own previous predictions as input, their prediction accuracy decreased, demonstrating the devastating effect of small compounding deviations propagated through time (Fig. 4C).
@@ 4 Discussion
In this paper, we presented theoretical results demonstrating that feed-forward and recurrent neural networks with fixed random weights but learnable biases can approximate arbitrary functions with high probability. We showcased the expressivity of bias-learned networks in auto-regressive modelling, multi-task learning, dynamic pattern generation, and dynamical system forecasting. Finally, we interrogated the representations learned by bias learning via analysis of task specialization, comparison with mask-learning, and an eigenvalue analysis in recurrent networks. In what follows we discuss key insights, limitations, and future directions.
Our results highlight three key insights. First, certain activation functions enable our universal approximation results when weights are drawn from a uniform distribution on a hyper-cube with any strictly positive edge length. Characterizing functions that do or do not support this property–we speculate that non-differentiable points might be an important component–and the link between hyper-cube edge length and network scaling (Remark 1 on Lemma B.4 in Appendix) both seem worthy of future work. Second, in the context of multi-tasking, we showed that bias learning finds solutions that rely on task-selective clusters (Fig.1), similar to the fully trained case Yang et al. (2019). Third, we believe our finding that bias learning yields solutions that are different, though related to, mask learning (Fig.2) suggests that further investigation of our method might shed light on this and other non-synaptic learning approaches.
@@
Our three main study limitations inspire directions for future research. First, the mathematical convergence results for dynamical systems are only point-wise over initial conditions, and are for finite-time trajectories. We believe that one might strengthen these results at least to $L^{p}$ convergence over trajectories, and that one might be able to break the finite-time limitations by studying convergence in stationary distribution. Second, our theory massively overestimates how wide the hidden layer of a bias-learned network should scale to reach a desired level of error (Remark 2 on Lemma B.4 in Appendix). This limitation might be addressed by drawing on scaling arguments from the SLTH literature (see e.g. Malach et al. (2020)). Third, a potential confounding factor in our comparison of bias and mask learning is that our mask learning approach used a learning schedule in the steepness parameter for the soft-masks. It is possible that the altered learning dynamics due to this scheduling contributed to mask and bias learning finding different solutions. Addressing this confound is an important direction for future work.
@@
Future directions should focus on capturing greater biological detail, and better hidden layer scaling. Experimental Ferguson and Cardin (2020) and theoretical work Wyrick and Mazzucato (2021); Ogawa et al. (2023) showed that neural pathways that modulated biases, like firing threshold or tonic inputs, may effect other neuronal properties, like neuron input-output gain. As our proofs rely on masking, they demonstrate universal approximation not just for bias learning but for any learned mechanism that can mask neurons, possibly including gain modulation. Exploring the flexibility of paradigms where gain (Stroud et al. (2018)) and biases are learned in concert is an interesting direction to exploit this. Moreover, the observed distribution of synaptic weights in the brain is not uniform but long-tailed Song et al. (2005), and searching for weight distributions that improve expressivity and hidden-layer scaling is an exciting future direction. If hidden-layer scaling can be improved, bias-learning might open up options for temporal credit assignment that perform poorly due to the $N^{2}$ scaling of synapses. Alternatively, it would be fascinating if, regardless of whether learned parameters are weights or biases, one needs roughly the same number of parameters to achieve a given degree of task performance.
Combined with decades old results on adaptive behaviour in networks without weight changes Cotter and Conwell (1990), and related findings in ML (see in-context learning Brown et al. (2020)), we hope that this study will inspire more research on learning phenomena that transcend synaptic adaptation.
Code Availability
Code will be released upon publication.
@@ Acknowledgements
E.W. was supported by an NSERC CGS D scholarship, and wishes to thank the other members of the Lajoie lab for support and for helpful discussions. T.J. was supported by an NSERC CGS M scholarship. M.G.P. was supported by grant the Fonds de recherche du Québec – Santé (chercheurs-boursiers en intelligence artificielle). L.M. was partially supported by National Institutes of Health grants R01NS118461, R01MH127375 and R01DA055439 and National Science Foundation CAREER Award 2238247. GL acknowledges CIFAR and Canada chair programs
Appendix A Random Neural Network Formulation
@@
$\tilde{r}_{m}^{\mathcal{M}}=-\alpha r_{0}+\mathcal{M}\odot\beta\phi(Wr_{0}+Bx+b),\quad\tilde{y}_{m}^{\mathcal{M}}=A\tilde{r}_{m}^{\mathcal{M}},$ (A.1)
@@
Consider the case where the $i^{th}$ element of $r_{0}$ is 00 whenever $\mathcal{M}_{i}=0$. Then, regardless of whether Eq.A.1 represents a feed-forward network or the transition function for an RNN, the masked units will always be zero. We can thus simply track the $n$ units that correspond with $1$’s in $\mathcal{M}$ as the outputs, $y^{\mathcal{M}}$ will depend solely on these. We observe that the behaviour of these units can be described by the following network:
$r_{m}^{\mathcal{M}}=-\alpha r_{0}+\beta\phi(W^{\mathcal{M}}r_{0}+B^{\mathcal{M}}x+b^{\mathcal{M}}),\quad y_{m}^{\mathcal{M}}=A^{\mathcal{M}}r_{m}^{\mathcal{M}}.$ (A.2)
It is networks of the form of Eq.A.2 that will be the primary subject of study in what follows. Note that the $\sim$’, over the $r$, is dropped to denote the fact that $r$ is a different vector on account of dropping the zero units. In the feed-forward case we use subscripts, as we have done above, to denote hidden layer width. Whenever we discuss RNNs or dynamical systems we will instead use the subscript to denote time.
Appendix B Proofs from Section 2.1
@@
$y_{n}(\theta)=\alpha^{2}\sum_{i=1}^{n}\frac{A_{:i}}{\alpha}\phi\bigg{(}\frac{B_{i:}}{\alpha}x+\frac{b_{i}}{\alpha}\bigg{)}=\alpha^{2}y_{n}\Big{(}\frac{\theta}{\alpha}\Big{)}.$ (B.1)
@@
$y_{n}(\theta)=\sum_{i=1}^{\alpha^{2}n}\tilde{A}_{:i}\phi\big{(}\tilde{B}_{i:}x+\tilde{b}\big{)}=y_{\alpha^{2}n}(\tilde{\theta}),$ (B.2)
where $\tilde{\theta}=[\frac{\theta_{1}}{\alpha},\dots,\frac{\theta_{n}}{\alpha},\dots\frac{\theta_{1}}{\alpha},\dots,\frac{\theta_{n}}{\alpha}]$ so that each element is simply a re-scaled and repeated version of the original parameters; we have $\alpha^{2}$ repeats for each term to replace the $\alpha^{2}$ factor in the LHS of Eq. B.1.
Now, given an arbitrary compact set $U\in\mathbb{R}^{d}$, continuous function $h:\mathbb{R}^{d}\to\mathbb{R}^{l}$, and $\varepsilon>0$, by the universal approximator theory (see e.g. Leshno et al. [1993], Hornik [1993]) we can find $n$ such that
$\int_{U}||h(x)-y_{n}(x,\theta)||\leq\epsilon$ (B.3)
holds. Because $n$ is finite we can bound every individual (scalar) parameter by $M$, for some sufficiently large $M$. Suppose we want the parameters to be bounded instead by $\gamma$ with $M>\gamma>0$. If we select $\alpha\in\mathbb{N}$ s.t. $\alpha>\frac{M}{\gamma}$ then we can find $y_{\alpha^{2}n}(x,\tilde{\theta})$ such that $y_{\alpha^{2}n}(x,\tilde{\theta})=y_{n}(x,\theta)$. Thus we have found a parameter-bounding ReLU neural network satisfying Eq.B.3, completing the proof.
@@
The following lemma constitutes the core of Theorem 1. It shows that one can achieve universal approximation, in the sense needed for the theorem, using masking. The theorem then follows by manipulating biases to achieve masking. As mentioned in the main text, the following lemma is very closely related to past results on the SLTH over units for MLPs with one hidden layer. Specifically, we believe that one should be able to prove an analogue to our theorem by combining Theorem 3.2 in Malach et al. [2020] with Theorem 1 of Rahimi and Recht [2008] (or a similar result on learning with random networks). We leave the details of this to future studies.
@@
First, we find a neural network with parameters that approximate the desired function $h$. Given the assumptions on $\phi$, we can use Proposition D.1 to find $n$ and parameters $\theta^{*}=\{A^{*},B^{*},b^{*}\}$ such that
[…]
@@
[…]
@@
$|\theta^{*}_{i}-\theta_{i}^{\mathcal{M}}|<\varepsilon=\frac{\epsilon}{2|U|\max\big{(}nl(K_{\phi}\bar{\gamma}[1+M_{x}]+M_{\phi}),1\big{)}}$ (B.6)
@@
$||y_{m}^{\mathcal{M}}(x)-y_{n}(x)||\leq nl[K_{\phi}\bar{\gamma}(1+M_{x})+M_{\phi}]\varepsilon\leq\frac{\epsilon}{2|U|},$ (B.7)
where, in addition to Eq.B.6, we use the assumptions on $\phi$ and $U$ stated before the start of the proof and repeated application of the triangle inequality (see Lemma 4 for the derivation of the above bound).
Integrating the error over the domain and using Eq.B.5 and Eq.B.7 gives
[…]
with probability $1-\delta$.
@@
Observe that, once we have choosen an $m$ satisfying the desiderata of Lemma B.4, because $\phi$ is assumed to be $\gamma$-bias-learning, $m$ is some finite value and all variables that make up the input of $\phi$ are bounded, we can implement the mask by setting $b_{i}$ to be very negative for every $i$ such that $\mathcal{M}_{i}=0$. For every $b_{i}$ such that $\mathcal{M}=1$ we simply leave $b_{i}$ at its original randomly chosen value. ∎
@@
Assume $d=l$, that is, the output and input spaces are the same. Then the results of Lemma B.4 and Theorem 1 also hold for res-nets; that is, networks whose output is of the form $x+y_{m}^{\mathcal{M}}(x)$.
@@
This follows by observing that $h(x)+x$ is also a continuous function and then replacing $h(x)$ with $h(x)+x$ in Eq.B.4 and rearranging. ∎
Remark: While the error can be made arbitrarily small, the limit of zero error itself is undefined. This is because our proof relies on first approximating the given smooth function with a neural network with all parameters tuned and then approximating this second network using bias-learning to pick-out a matching sub-network from a large random reservoir; the probability of perfectly matching the fully tuned network with the bias-learned network is zero. This could be addressed by using an integral representation for continuous functions instead of directly using a finite-width neural network to approximate the given function (see e.g. Rahimi and Recht [2008], Li et al. [2023]). As one will see below, this remark also applies to the recurrent neural network result.
Appendix C Proof from Section 2.2
Analogous to the section containing the feed-forward proofs, we first state and prove a lemma which comprises the core of the proof for recurrent neural networks. This lemma shows that one can achieve universal approximation with high probability using masking in a randomly initialized RNN, and in this way provides a proof of the SLTH over units for RNNs. The proof of the main theorem in this section then follows quite straightforwardly.
@@
Consider a discrete time, partially observed dynamical system of the form of, and satisfying the same conditions as, the one in Eq.3. Let $0<T<\infty$, initial condition $z_{0}\in U_{z}$, input $x_{t}\in U_{x}$ $\forall t$, $\epsilon>0$ and $\delta\in(0,1)$. Then we can find an RNN initial condition $r_{0}$ and a layer width $m\in\mathbb{N}$ such that with probability at least $1-\delta$$\exists\mathcal{M}\in\{0,1\}^{m}$ satisfying the following:
$\sum_{t=1}^{T}||y_{t}-y_{t}^{\mathcal{M}}||<\epsilon.$ (C.1)

### v3

@@ 3.4 Bias learning non-autonomous dynamical systems with RNNs
The performance of the bias-learned RNNs scaled, as a function of trainable model parameters, in a qualitatively similar fashion to the bias-trained FNNs (Fig. 4A; see Fig. E.2D for scaling as function of layer width). Both the fully-trained and sufficiently wide bias-learned networks accurately predicted future points of the Lorenz system, evidenced by a consistent $R^{2}$ metric of $>0.99$ ($n$=5) achieved by networks with a hidden-layer width of 1024 on a window of the Lorenz time-series held out during training (Fig. 4B). However, when the networks were fed their own previous predictions as input, in place of the ground-truth time series, in an autoregressive, ‘self-sustained’ fashion, their prediction accuracy decreased, demonstrating the damaging effect of small compounding deviations propagated through time (Fig. 4C).
3.5 Bias-learned Motor control with RNNs
Finally, we tested whether bias learning could solve a center-out reaching task, a paradigm routinely used to study motor control in human and non-human primates Ashe & Georgopoulos (1994); Shadmehr & Mussa-Ivaldi (1994). Starting from the center of the workspace, the subject must move the selected end effector (e.g., their right hand) to reach several peripheral targets placed equidistantly on a circle. We modelled this task using an RNN with random weights and learned biases, where linear readouts controlled a point-mass arm (a unit mass modeling the arm’s behavior; see Methods). The task objective was to reach the peripheral targets in 1 second, with near-zero velocity and force at the end of movement, which were imposed using regularization terms in the loss function. A 1,024-units network required approximately $10^{4}$ training epochs—where one epoch involved presenting all targets once—to solve the task (Fig. 5A). (Note that a parameter-matched, fully-trained RNN can also solve this task, in $\sim 10^{3}$ epochs (Fig. E.5).) Importantly, a single set of biases was used to reach all 6 targets for each trained network. The resulting trajectories successfully reached the targets (Fig. 5B, dark curves and black targets), and exhibited the characteristic bell-shaped velocity profile Harris & Wolpert (1998) (Fig. 5C, top). Crucially, the trained network generalized decently to new targets on the circle (Fig. 5B, light curves and grey targets; Fig. 5C, bottom). To achieve this, the network had to produce both acceleration and deceleration when given information about the Cartesian position of a target never seen in the training period. These results highlight the flexibility of bias learning in generating diverse open-loop controls.
Figure 5: Center-out reaching task. A. Training loss for 3 network initializations. B. Trajectories for the trained (black) and tested (grey) targets. C. Speeds ($(\dot{x}^{2}+\dot{y}^{2})^{1/2}$) for the trained (top) and tested (bottom) targets (mean $\pm$ SD across targets).
@@ 4 Discussion
In this paper, we presented theoretical results demonstrating that FNNs and RNNs with fixed random weights but learnable biases can approximate arbitrary functions with high probability. We showcased the expressivity of bias-learned networks in multi-task, times-series, and motor control tasks, and interrogated their learned representations by analyzing task selectivity and comparing with mask-learning in FNNs, and by analyzing eigenvalues in RNNs.
@@
We underline four limitations of our study that might inspire future research. First, the convergence results for dynamical systems were only for finite-time trajectories. One could overcome this limitation by studying convergence in stationary distribution. Second, a potential confounding factor in our comparison of bias and mask learning is that the latter approach used a learning schedule in the steepness parameter for the soft masks. It is possible that the altered learning dynamics due to this scheduling contributed to mask and bias learning finding different solutions. Addressing this confound is an important direction for future work. Third, a better grasp of bias learning scaling–how layer width or parameter count increases as a function of performance–is needed. Our theory does not shed light here as it significantly overestimates how a bias-learned network should scale (Remark 2 on Lemma 3 in Appendix). Some insight can be gleaned by noting that, with bias-learning activations, tuned biases can express any mask-learned solution. Thus, results showing that mask learning layer widths scale polynomially in the inverse error and the size of a performance-matched random feature model (see Malach et al. (2020) Theorem 3.2) represent a worst-case scaling for bias learning. Whether bias learning scales better than mask learning could be a good starting point for addressing scaling. Fourth, while our results appeared to generalize beyond one-layer networks (Fig. E.2E), a more detailed study of bias-learned deep FNNs is left to future work.
@@
We lastly highlight the need for greater biological detail in future studies of bias learning. Past experimental Ferguson & Cardin (2020) and theoretical work Wyrick & Mazzucato (2021); Ogawa et al. (2023) showed that neural mechanisms modulating biases, like firing threshold or tonic inputs, may effect other neuronal properties, like neuron input-output gain. As our proofs rely on masking, they demonstrate universal approximation not just for bias learning but for any learned mechanism that can mask neurons. Exploring paradigms where gain (Stroud et al. (2018)) and biases are learned in concert could be an interesting direction to exploit this. Finally, the observed distribution of synaptic weights in the brain is not uniform but long-tailed Song et al. (2005), highlighting the need for bias learning models with more structured weight initializations. We hypothesize that structure in the weights intermediate between fully random and fully learned might yield an optimal combination of performance and training efficiency. If such structure improves hidden-layer scaling, this could enable bias learning to support temporal credit assignment algorithms that struggle in weight space due to the $N^{2}$ scaling of synapses. A fascinating alternative is that the same number of parameters could achieve a given task performance, regardless of whether they are weights or biases.
@@ Acknowledgements
E.W. was supported by an NSERC CGS D scholarship, and wishes to thank the other members of the Lajoie lab for support and for helpful discussions. T.J. was supported by an NSERC CGS M scholarship. M.G.P. was supported by grant the Fonds de recherche du Québec – Santé (chercheurs-boursiers en intelligence artificielle). L.M. was partially supported by National Institutes of Health grants R01NS118461, R01MH127375 and R01DA055439 and National Science Foundation CAREER Award 2238247. GL acknowledges CIFAR and Canada chair program.
Appendix A Biological Bias-Related Mechanisms
System Bias model Mechanism Effect Refs
Working memory Dopamine projections to cortical circuit Bias of NMDA or GABA conductances Enhance memory signal-to-noise ratio and resistance to distractors Brunel & Wang (2001)
Behavioral modulation of visual responses Thalamic projections to visual cortex Bias to pyramidal and somatostatin cells Paradoxical reversal of somatostatin cell responses to running Garcia del Molino et al. (2017)
Motor sequences Basal ganglia projections to thalamus Inhibitory biases of thalamic neurons Toggle on/off different motor primitives Logiaco et al. (2021)
Motor sequences Secondary motor cortex projections to primary motor cortex (M1) Biases of M1 neurons Select initial conditions (anticipatory activity) for motor primitives Recanatesi et al. (2022); Mazzucato (2022)
Short-term motor adaptation Premotor cortex projections to primary motor cortex (M1) Recruitment of M1 neurons Shift pre-movement preparatory states Perich et al. (2018); Feulner et al. (2022)
Expectation and taste processing Amygdalar projections to gustatory cortex Biases of cortical neurons Acceleration of taste coding via gain modulation Mazzucato et al. (2019); Vincis & Fontanini (2016); Haley et al. (2020)
Movements and visual processing Thalamic projections to visual cortex Biases of cortical neurons Acceleration of visual coding via gain modulation Wyrick & Mazzucato (2021)
Arousal and auditory processing Neuromodulatory projections to auditory cortex Biases of cortical neurons Optimal encoding of auditory stimuli at intermediate arousal Papadopoulos et al. (2024)
Table 1: Potential biological mechanisms for bias-related plasticity in the brain.
Appendix B Mathematical Proofs
Throughout the appendix the proofs are restated for ease of reference. We will always take |||||$\cdot$||| | ⋅ | | to be the $1$-norm unless stated otherwise.
B.1 Random Neural Network Formulation
@@
$\tilde{r}_{m}^{\mathcal{M}}=-\alpha r_{0}+\mathcal{M}\odot\beta\phi(Wr_{0}+Bx+b),\quad\tilde{y}_{m}^{\mathcal{M}}=A\tilde{r}_{m}^{\mathcal{M}},$ (4)
@@
Consider the case where the $i^{th}$ element of $r_{0}$ is 00 whenever $\mathcal{M}_{i}=0$. Then, regardless of whether Eq.4 represents a feed-forward network or the transition function for an RNN, the masked units will always be zero. We can thus simply track the $n$ units that correspond with $1$’s in $\mathcal{M}$ as the outputs, $y^{\mathcal{M}}$ will depend solely on these. We observe that the behaviour of these units can be described by the following network:
$r_{m}^{\mathcal{M}}=-\alpha r_{0}+\beta\phi(W^{\mathcal{M}}r_{0}+B^{\mathcal{M}}x+b^{\mathcal{M}}),\quad y_{m}^{\mathcal{M}}=A^{\mathcal{M}}r_{m}^{\mathcal{M}}.$ (5)
It is networks of the form of Eq.5 that will be the primary subject of study in what follows. Note that the $\sim$’, over the $r$, is dropped to denote the fact that $r$ is a different vector on account of dropping the zero units. In the feed-forward case we use subscripts, as we have done above, to denote hidden layer width. Whenever we discuss RNNs or dynamical systems we will instead use the subscript to denote time.
B.2 Proofs from Section 2.1
@@
$y_{n}(\theta)=\alpha^{2}\sum_{i=1}^{n}\frac{A_{:i}}{\alpha}\phi\bigg{(}\frac{B_{i:}}{\alpha}x+\frac{b_{i}}{\alpha}\bigg{)}=\alpha^{2}y_{n}\Big{(}\frac{\theta}{\alpha}\Big{)}.$ (6)
@@
$y_{n}(\theta)=\sum_{i=1}^{\alpha^{2}n}\tilde{A}_{:i}\phi\big{(}\tilde{B}_{i:}x+\tilde{b}_{i}\big{)}=y_{\alpha^{2}n}(\tilde{\theta}),$ (7)
where $\tilde{\theta}=[\frac{\theta_{1}}{\alpha},\dots,\frac{\theta_{n}}{\alpha},\dots\frac{\theta_{1}}{\alpha},\dots,\frac{\theta_{n}}{\alpha}]$ so that each element is simply a re-scaled and repeated version of the original parameters; we have $\alpha^{2}$ repeats for each term to replace the $\alpha^{2}$ factor in the LHS of Eq. 6.
Now, given an arbitrary compact set $U\in\mathbb{R}^{d}$, continuous function $h:\mathbb{R}^{d}\to\mathbb{R}^{l}$, and $\varepsilon>0$, by the universal approximator theory (see e.g. Leshno et al. (1993); Hornik (1993)) we can find $n$ such that
$\sup_{x\in U}||h(x)-y_{n}(x,\theta)||\leq\epsilon$ (8)
holds. Because $n$ is finite we can bound every individual (scalar) parameter by $M$, for some sufficiently large $M$. Suppose we want the parameters to be bounded instead by $\gamma$ with $M>\gamma>0$. If we select $\alpha\in\mathbb{N}$ s.t. $\alpha>\frac{M}{\gamma}$ then we can find $y_{\alpha^{2}n}(x,\tilde{\theta})$ such that $y_{\alpha^{2}n}(x,\tilde{\theta})=y_{n}(x,\theta)$. Thus we have found a parameter-bounding ReLU neural network satisfying Eq.8, completing the proof.
@@
The following lemma constitutes the core of Theorem 1. It shows that one can achieve universal approximation, in the sense needed for the theorem, using masking. The theorem then follows by manipulating biases to achieve masking.
@@
First, we find a neural network with parameters that approximate the desired function $h$. Given the assumptions on $\phi$, we can use Proposition 19 to find $n$ and parameters $\theta^{*}=\{A^{*},B^{*},b^{*}\}$ such that
$\sup_{x\in U}||h(x)-y_{n}(x,\theta^{*})||\leq\frac{\epsilon}{2},$ (10)
@@
We now make two observations: first, all choices of $x$ are from a compact set, $U$, by assumption and the parameters of a given random or non-random neural net are also from a compact set–the $n$ dimensional hyper-cube with edge length $2\bar{\gamma}$. Second, the function $y_{n}$ is a continuous function of $x$ and the parameters. By these two observations the function $y_{n}$ is a continuous function on the compact product space of inputs and parameters, and thus admits a Lipshitz constant, $K_{n}$. This will come in handy momentarily.
@@
$|\theta^{*}_{i}-\theta_{i}^{\mathcal{M}}|<\varepsilon=\frac{\epsilon}{2K_{n}n(d+l+1)}$ (11)
@@
$||y_{m}^{\mathcal{M}}(x)-y_{n}(x,\theta^{*})||\leq K_{n}\big{(}||\theta^{\mathcal{M}}-\theta^{*}||+||x-x||\big{)}\leq K_{n}n(d+l+1)\varepsilon\leq\frac{\epsilon}{2},$ (12)
where, in addition to Eq.11, we used the fact that $y_{m}^{\mathcal{M}}(x)=y_{n}(x,\theta^{\mathcal{M}})$, the continuity and compactness mentioned above, and repeated application of the triangle inequality. Importantly, this bound holds for all $x\in U$. Because $\phi$ is assumed continuous, the function $f(x)=||y_{m}^{\mathcal{M}}(x)-y_{n}(x)||$ is also continuous. By the extreme value theorem $\exists$ $x^{\star}\in U$ such that $\sup_{x\in U}f(x)=f(x^{\star})$. Since $x^{\star}\in U$ the bound from Eq.12 applies and we have:
$\sup_{x\in U}||y_{m}^{\mathcal{M}}(x)-y_{n}(x)||=||y_{m}^{\mathcal{M}}(x^{\star})-y_{n}(x^{\star})||\leq\frac{\epsilon}{2}.$ (13)
Using the triangle inequality, Eq.10, and Eq.13 gives $\sup_{x\in U}||h(x)-y_{m}^{\mathcal{M}}(x)||\leq\epsilon$ with probability $1-\delta$.
Connections with Malach et al. 2020: As mentioned in the main text, the previous lemma is closely related to past results on the SLTH over units for MLPs with one hidden layer. In Theorem 3.2 of Malach et al. (2020), it is proven that one can match the performance of a random feature model (an MLP where only the output layer is trained) by scaling the output of a mask-learned network. This theorem differs from ours in three important ways. First, it proves a result for a potentially different class of activation functions, second, it requires a scaling of the output of the network–unlike our proof which does not require this–and, third, it compares the performance of mask-learned networks to random feature models, rather than directly proving that mask-learned networks can approximate wide classes of functions. We believe that one should be able to achieve a result similar to ours by combining Malach et al.’s Theorem 3.2 with Theorem 1 of Rahimi & Recht (2008) (or a similar result on learning with random networks), but we leave the details of this to future studies.
@@
Observe that, once we have choosen an $m$ satisfying the desiderata of Lemma 9, because $\phi$ is assumed to be $\gamma$-bias-learning, $m$ is some finite value and all variables that make up the input of $\phi$ are bounded, we can implement the mask by setting $b_{i}$ to be very negative for every $i$ such that $\mathcal{M}_{i}=0$. For every $b_{i}$ such that $\mathcal{M}=1$ we simply leave $b_{i}$ at its original randomly chosen value. ∎
@@
Assume $d=l$, that is, the output and input spaces are the same. Then the results of Lemma 9 and Theorem 1 also hold for res-nets; that is, networks whose output is of the form $x+y_{m}^{\mathcal{M}}(x)$.
@@
This follows by observing that $h(x)+x$ is also a continuous function and then replacing $h(x)$ with $h(x)+x$ in Eq.9 and rearranging. ∎
Remark: While the error can be made arbitrarily small, the limit of zero error itself is undefined. This is because our proof relies on first approximating the given smooth function with a neural network with all parameters tuned and then approximating this second network using bias-learning to pick-out a matching sub-network from a large random reservoir; the probability of perfectly matching the fully tuned network with the bias-learned network is zero. This could be addressed by using an integral representation for continuous functions instead of directly using a finite-width neural network to approximate the given function (see e.g. Rahimi & Recht (2008); Li et al. (2023)). As one will see below, this remark also applies to the recurrent neural network result.
B.3 Proof from Section 2.2
Analogous to the section containing the feed-forward proofs, we first state and prove a lemma which comprises the core of the proof for recurrent neural networks. This lemma shows that one can achieve universal approximation in the $L^{\infty}$ norm over trajectories sense (see section §2.2) with high probability using masking in a randomly initialized RNN, and in this way provides a proof of the SLTH over units for RNNs. The proof of the main theorem in this section then follows quite straightforwardly.
@@
Consider a discrete time, partially observed dynamical system of the form of, and satisfying the same conditions as, the one in Eq.3. Let $0<T<\infty$, $x_{t}\in U_{x}$ $\forall t$, $\epsilon>0$ and $\delta\in(0,1)$. Then we can find an RNN with appropriately chosen initial conditions and a layer width $m\in\mathbb{N}$ such that with probability at least $1-\delta$$\exists\mathcal{M}\in\{0,1\}^{m}$ satisfying the following:
$\sup_{z_{0}\in U_{r},x_{0:(T-1)}\in U_{\vec{x}}}\sum_{t=1}^{T}||y_{t}-y_{t}^{\mathcal{M}}||<\epsilon.$ (14)